# Google Research

### v1

@@ 4.1 Choosing the Set of Keywords $\mathcal{K}$
[…]
@@
We compute the document frequency (DF) of all terms with additive Laplace noise to satisfy DP. Using these noisy DF values with token frequency (TF), we select 3–5 keywords for each conversation. We then apply partition selection to privately identify significant keywords (Desfontaines et al., 2022), followed by LLM refinement to obtain $\sim$200 final keywords. This approach is better suited when direct LLM access to conversations raises additional privacy concerns.
@@
We first use an LLM to select 3–5 keywords for each conversation independently. This produces a large candidate set of keywords across all conversations. We then apply partition selection to privately identify the most representative keywords while satisfying $(\varepsilon,\delta)$-DP. This yields several thousand keywords, which we refine using the LLM to produce a final set of $\sim$200 keywords.
Iterative TF-IDF Refinement (KwSet-Public).
As a non-private baseline, we first use TF-IDF to extract an initial set of keywords from a subset of the corpus. We then sequentially update this set using an LLM. In each iteration, we present the LLM with the current set and new conversations, prompting it to output words_to_remove and words_to_add. This sequential refinement continues until the keyword set stabilizes or reaches a desired size.
@@
In this hybrid approach, we leverage KwSet-Public while maintaining privacy guarantees. We provide the LLM with KwSet-Public along with several private conversations, asking it to output relevant keywords only when necessary. We then apply partition selection over these selected keywords to ensure DP. We propose this approach to address potential distribution shifts between the public conversations used in constructing KwSet-Public and the private conversations we aim to analyze. This hybrid method allows for adaptation to new topics or terms in the private dataset while still benefiting from the high-quality foundation of the public keyword set.
@@ 5 Evaluations
Our evaluation strategy adopts a multifaceted approach to assess the privacy-utility tradeoff in our DP pipeline. We compare summaries generated by our private pipeline against those from the public Clio pipeline Tamkin et al. (2024) on identical conversation sets, using the public outputs as a reasonable proxy for ground truth—an advantage not available in the original work.
We employ both non-LLM methods (comparing key phrases, n-grams, topics, and embedding similarities) and LLM-based evaluations (comparative quality assessments and independent scoring) to comprehensively measure how well our privacy-enhanced summaries preserve the essential information captured in the public summaries. This approach allows us to quantify both the utility preservation and information retention of our DP pipeline while maintaining stronger formal privacy guarantees than the original framework. Additionally, we evaluate privacy protection through AUC score comparisons between the public and private pipelines, providing a more complete picture of the privacy-utility tradeoff in our system, which will be presented in the later sections (See §E).
@@ 5.1 Evaluation Methodology
    Lexical Content Analysis: We extract key phrases, noun chunks, and TF-IDF keywords from both sets of summaries and compute similarity metrics including Jaccard similarity and cosine similarity between these feature sets.
@@
    N-gram and Topic Analysis: We analyze tokens and 2-grams extracted from both sets of summaries to assess content preservation at different granularities. In particular, we use BERTopic (Grootendorst, 2022) to extract topics from both sets of summaries and measure their overlap.
@@ 5.2 Experimental Setup & Results
We try to analyze the conversations from the popular public datasets LMSYS-1M-Chat (Zheng et al., 2023). The LLM we used is Gemini-2.0-flash-001. In constructing KwSet-Public, we use the public WildChat dataset (Zhao et al., 2024). In particular, we set the cluster size privacy parameter $\varepsilon_{\text{size}}=1$ for cluster size verification and implement a 1-DP keyword set generation.
We evaluate our approach across different privacy parameter settings and keyword set configurations. We consider varying values of $\varepsilon_{\mathrm{c}}$ and $\varepsilon_{\mathrm{hist}}$, while keeping the keyword set to be generated using KwSet-TFIDF. We also fix $\varepsilon_{\mathrm{c}}=8.0$ and $\varepsilon_{\mathrm{hist}}=4.0$ and evaluate with different choices of keyword sets.
@@
Similarity lower is better) (DP preferred)
@@ 5.3 Discussion
We have an LLM evaluator independently score private and public summaries on a scale from 1-5 (where 1 means very poor and 5 means excellent) based on how well they summarize the original conversations. The private summaries are slightly better than public summaries, but both achieve very low scores on average ($<$1.4). This suggests that there might be significant room for improvement in the overall quality of conversation summarization, regardless of privacy considerations, and highlights the challenging nature of the summarization task itself.
These findings highlight both the capabilities and limitations of DP text summarization. Our framework demonstrates that it is possible to generate meaningful summaries while providing formal privacy guarantees, but practitioners should carefully consider the privacy-utility trade-off when configuring the system for real-world applications. However, qualitative analysis of specific examples (see Appendix C) reveals important limitations in our private summaries that may not be captured by automated evaluation metrics.
@@ 6 Discussion and Future Work
Our current pipeline exclusively addresses differential privacy (DP) at the record-level, where privacy guarantees are provided when only a single conversation is modified. However, in practical scenarios, individual users typically contribute multiple conversations with large language models (LLMs). This motivates an important open research problem: extending the privacy notion to user-level differential privacy, wherein the privacy guarantees hold when all conversations from a single user are removed.
@@
While our work focuses on formal differential privacy guarantees for released summaries, we acknowledge that comprehensive privacy protection requires addressing multiple risk vectors. Our approach provides one layer of protection against information leakage through published analyses, but does not address other important concerns such as centralized data collection risks, third-party model processing vulnerabilities, or inference attacks on the underlying conversation data. Future work should explore how formal DP guarantees can be integrated with other privacy-preserving techniques (such as federated learning, secure multi-party computation, or local differential privacy) to create more comprehensive privacy frameworks for conversation analysis systems.
@@ Appendix A Formal Details of DP Tools
  • [Parallel Composition] If $\mathcal{M}:\Omega^{*}\to\mathcal{R}$ satisfies $(\varepsilon,\delta)$-DP, then for any disjoint collection of subsets $\Omega_{1},\Omega_{2},\ldots,\Omega_{r}\subseteq\Omega$, given a dataset $D\subseteq\Omega$, the mechanism that returns $(\mathcal{M}(D\cap\Omega_{1}),\mathcal{M}(D\cap\Omega_{2}),\ldots,\mathcal{M}(D\cap\Omega_{r}))$ also satisfies $(\varepsilon,\delta)$-DP.
@@ A.2 Private Partition Selection
The mechanism $\texttt{PPS}_{k,\varepsilon,\delta}(h_{1},\ldots,h_{n})$ returns a subset of elements for which the noisy count, after adding truncated discrete Laplace noise $\mathsf{DLap}_{\tau}(\varepsilon/k)$ is more than $\tau$.
@@
Output: The set $R\subseteq\bigcup_{i}h_{i}$.
$\varepsilon^{\prime},\delta^{\prime}\leftarrow\varepsilon/k,\delta/k$
@@
       $c_{\omega}\leftarrow|\{i~{}:~{}h_{i}\ni\omega\}|$ $\zeta_{\omega}\sim\mathsf{DLap}_{\tau}(\varepsilon)$
      if $c_{\omega}+z_{\omega}>\tau$ then
@@
Algorithm 3 Private Partition Selection $\texttt{PPS}_{\varepsilon,\delta}(x_{1},\dots,x_{n})$
@@ Appendix B Description of Simple-Clio
$\bm{e}\leftarrow\textsc{ExtractEmbeddings}(x_{i}):i\in[n]\}$
@@
       $s_{j}\leftarrow\textsc{LLM-Summarize}(\mathbf{samples}_{j})$ ;
@@
Appendix C Summary Comparison Examples

### v2

@@ 4.1 Choosing the Set of Keywords $\mathcal{K}$
The effectiveness of Urania crucially depends on the quality of the keyword set $\mathcal{K}$ used during cluster summarization. We explore four approaches for constructing keyword sets $\mathcal{K}$, with varying privacy guarantees: namely, KwSet-TFIDF and KwSet-LLM satisfy DP, KwSet-Public uses a public dataset to obtain keywords, and KwSet-Hybrid uses a hybrid approach. We describe these methods informally below and formally in §C. We discuss some related work in §E.
@@
We compute the document frequency (DF) of all terms with additive Laplace noise to satisfy DP. Using these noisy DF values with token frequency (TF), we select 3–5 keywords for each conversation. We then apply partition selection (Theorem 2.3) to privately identify significant keywords, followed by LLM refinement to obtain $\sim$200 final keywords (see §G.5 for prompt used for refining). This approach is better suited when direct LLM access to conversations raises additional privacy concerns.
@@
We first use an LLM to select 3–5 keywords for each conversation independently. This produces a large candidate set of keywords across all conversations. We then apply partition selection to privately identify the most representative keywords while satisfying $(\varepsilon,\delta)$-DP. This yields several thousand keywords, which we refine using the LLM to produce a final set of $\sim$200 keywords. See detailed prompts in §G.5.
Iterative NLP-based Refinement (KwSet-Public).
As a non-private baseline, we first use multiple NLP techniques (Named Entity Recognition, noun chunk extraction, and RAKE keyword extraction) to extract an initial set of keywords from a standalone (but related) public dataset. We then sequentially update this set using an LLM. In each iteration, we present the LLM with the current set of keywords and each new conversation, prompting it to output words_to_remove and words_to_add. This sequential refinement continues until the keyword set stabilizes or reaches a desired size; see detailed prompt in §G.6.
@@
In this hybrid approach, we leverage KwSet-Public while maintaining privacy guarantees. We provide the LLM with KwSet-Public along with several private conversations, asking it to output relevant keywords only when necessary. We then apply partition selection over these selected keywords to ensure DP. We propose this approach to address potential distribution shifts between the public conversations used in constructing KwSet-Public and the private conversations we aim to analyze. This hybrid method allows for adaptation to new topics or terms in the private dataset while still benefiting from the high-quality foundation of the public keyword set. See detailed prompt in §G.7.
@@ 5 Evaluations
Our evaluation strategy adopts a multifaceted approach to assess the privacy-utility tradeoff in our DP pipeline. We compare summaries generated by our private Urania pipeline against those from the (non-private) Simple-Clio on identical conversation sets, using the latter as a reasonable proxy for ground truth.
We employ both non-LLM methods (comparing key phrases, n-grams, topics, and embedding similarities) and LLM-based evaluations (comparative quality assessments and independent scoring) to comprehensively measure how well our privacy-enhanced summaries preserve the essential information captured in the public summaries. This approach allows us to quantify both the utility preservation and information retention of our DP pipeline while maintaining stronger formal privacy guarantees than the original framework. Additionally, we evaluate privacy protection through AUC score comparisons between the public and private pipelines, providing a more complete picture of the privacy-utility tradeoff in our system, which will be presented in the later sections (See §F).
@@ 5.1 Evaluation Methodology
    Lexical Content Analysis: We extract key phrases, noun chunks, and TF-IDF keywords from both sets of summaries and compute similarity metrics with Jaccard similarity between these feature sets.
@@
    N-gram and Topic Analysis: We analyze tokens and 2-grams extracted from both sets of summaries to assess content preservation at different granularities. Additionally, we use BERTopic (Grootendorst, 2022) to extract topics from both sets of summaries and measure their overlap.
@@
Finally, we propose a simple empirical privacy evaluation method using a membership inference-style attack to conceptually demonstrate that Urania is indeed more robust than Simple-Clio; details in §F.
@@ 5.2 Experimental Setup & Results
Similarity lower is better) (% DP preferred)
@@
We analyze the conversations from the popular public datasets LMSYS-1M-Chat (Zheng et al., 2023). Specifically we use gemini-2.0-flash-001 language model. In constructing KwSet-Public, we use the public WildChat dataset (Zhao et al., 2024). In particular, we set the cluster size privacy parameter $\varepsilon_{\text{size}}=1$ for cluster size verification and implement a 1-DP keyword set generation.
We evaluate our approach across different privacy parameter settings and keyword set configurations. We consider varying values of $\varepsilon_{\mathrm{c}}$ and $\varepsilon_{\mathrm{hist}}$, while keeping the keyword set to be generated using KwSet-TFIDF. We also fix $\varepsilon_{\mathrm{c}}=8.0$ and $\varepsilon_{\mathrm{hist}}=4.0$ and evaluate with different choices of keyword sets. Note that Urania with $\varepsilon=\infty$ is still distinct from Simple-Clio because the former generates cluster summaries via keywords, whereas the latter does so directly from individual summaries.
@@ 5.3 Discussion
We have an LLM evaluator independently score private and public summaries on a scale from 1-5 (where 1 means very poor and 5 means excellent) based on how well they summarize the original conversations. The private summaries are slightly better than public summaries, but both achieve very low scores on average (<<1.4). This suggests that there might be significant room for improvement in the overall quality of conversation summarization, regardless of privacy considerations, and highlights the challenging nature of the summarization task itself.
These findings highlight both the capabilities and limitations of DP text summarization. Our framework demonstrates that it is possible to generate meaningful summaries while providing formal privacy guarantees, but practitioners should carefully consider the privacy-utility trade-off when configuring the system for real-world applications. However, qualitative analysis of specific examples (see §D) reveals important limitations in our private summaries that may not be captured by automated evaluation metrics.
@@ 6 Discussion and Future Work
Our current pipeline exclusively addresses DP at the record-level, where privacy guarantees are provided when only a single conversation is modified. However, in practical scenarios, individual users typically contribute multiple conversations with large language models (LLMs). This motivates an important open research problem: extending the privacy notion to user-level DP, wherein the privacy guarantees hold when all conversations from a single user are removed.
@@
While our work focuses on formal DP guarantees for released summaries, we acknowledge that comprehensive privacy protection requires addressing multiple risk vectors. Our approach provides one layer of protection against information leakage through published analyses, but does not address other important concerns such as centralized data collection risks, third-party model processing vulnerabilities, or inference attacks on the underlying conversation data. Future work should explore how formal DP guarantees can be integrated with other privacy-preserving techniques (such as federated learning, secure multi-party computation, or local DP) to create more comprehensive privacy frameworks for conversation analysis systems.
@@ Appendix A Formal Details of DP Tools
  • [Parallel Composition] If $\mathcal{M}:\Omega^{*}\to\mathcal{R}$ satisfies $(\varepsilon,\delta)$-DP, then for any disjoint collection of subsets $\Omega_{1},\ldots,\Omega_{r}\subseteq\Omega$, given a dataset $D\subseteq\Omega$, the mechanism that returns $(\mathcal{M}(D\cap\Omega_{1}),\ldots,\mathcal{M}(D\cap\Omega_{r}))$ also satisfies $(\varepsilon,\delta)$-DP.
@@ A.2 Private Partition Selection
The mechanism $\texttt{PPS}_{k,\varepsilon,\delta}(h_{1},\ldots,h_{n})$ returns a subset of elements for which the noisy count, after adding truncated discrete Laplace noise $\mathsf{DLap}_{\tau}(\varepsilon/k)$ is more than $\tau$. The result for $k=1$ was provided by Desfontaines et al. (2022) and the result for general $k$ can be derived using the group privacy property of DP (see Lemma 2.2 in Vadhan (2017)).
@@
Output: A set $R\subseteq\bigcup_{i}h_{i}$.
[…]
@@
 $c_{\omega}\leftarrow|\{i~:~h_{i}\ni\omega\}|$
 $\zeta_{\omega}\sim\mathsf{DLap}_{\tau}(\varepsilon)$
 if $c_{\omega}+\zeta_{\omega}>\tau$ then
@@
Algorithm 3 Private Partition Selection $\texttt{PPS}_{k,\varepsilon,\delta}(x_{1},\dots,x_{n})$
@@ Appendix B Description of Simple-Clio
We formally describe the Simple-Clio algorithm in Algorithm 4 complementing the description in §3.
@@
$\bm{e}\leftarrow\{\textsf{ExtractEmbeddings}(x_{i}):i\in[n]\}$
@@
 $s_{j}\leftarrow\textsf{LLM-Summarize}(\mathbf{samples}_{j})$ ;
@@
Algorithm 4 $\textsc{Simple}\text{-}\textsc{Clio}_{k}$ : Non-Differentially Private Text Corpora Summarization
The sub-routines used in the above algorithm operate as follows:
  • $\textsf{ExtractEmbeddings}(x)$ uses an LLM to create a “summary” for $x$ (see prompt in §G.1), and thereafter uses an embedding model that maps the summary to a real-valued vector. Specifically, we use all-mpnet-base-v2 embedding model (Reimers & Gurevych, 2019; 2022).
  • $\textsc{KMeans}_{k}(\bm{e})$ applies the $k$-means algorithm that partitions the vectors $\bm{e}$ into $k$ clusters that minimizes the $k$-means objective.
  • SampleConversations draws a small (e.g., $10$) number of random conversations from the specified cluster.
  • LLM-Summarize uses an LLM to generate a summary associated to the cluster using the sampled conversations from the cluster (see prompt in §G.2).
Appendix C Keyword Selection Algorithms
We formally describe the methods for extract keywords from documents within each cluster, complementing the description in §4.1. In these methods, we apply the extraction on the concatenation of the various facets of the inputs $x_{i}$’s. However, for simplicity, we write the pseudocode assuming they are applied on the raw inputs.
TF-IDF-Based Selection with Partition Selection (KwSet-TFIDF).
This method proceeds by constructing a token frequency matrix $\mathbf{TF}\in\mathbb{Z}^{n\times|V|}$ where $\mathbf{TF}_{i,j}$ is the number of times the token $j$ appears in input $x_{i}$. We normalize the token frequency matrix such that each column has $\ell_{1}$ norm at most $w_{\max}$.
Additionally, we construct a document frequency vector $\mathbf{df}\in\mathbb{Z}^{|V|}$, where $\mathbf{df}_{j}$ is the number of times the token $j$ appears across all documents. This is privatized to obtain $\widetilde{\mathbf{df}}$ by adding discrete Laplace noise. And similarly, the number of documents is estimated as $\tilde{n}$ by adding discrete Laplace noise to $n$.
@@ Appendix C Keyword Selection Algorithms
Parameters: Maximum keywords per conversation $k$, maximum document weight $w_{\max}$, maximum total keywords $K$.
Parameters: Privacy parameters $\varepsilon_{\mathrm{idf}},\varepsilon_{\mathrm{sel}}>0$, $\delta_{\mathrm{sel}}\in[0,1]$.
Input: Input dataset $\bm{x}=(x_{1},x_{2},\ldots,x_{n})$ of structured facets
Output: Differentially private keyword set $\mathcal{K}$
1ex
// Step 1: Build TF matrix and apply L1 clipping
$\mathbf{TF},\mathcal{V}\leftarrow\textsf{CountVectorizer}(\bm{x})$ ;
// Build term frequency matrix and token list
for $i\leftarrow 1$ to $n$ do
 if $\|\mathbf{TF}_{i,:}\|_{1}>w_{\max}$ then
    $\mathbf{TF}_{i,:}\leftarrow\mathbf{TF}_{i,:}\cdot\frac{w_{\max}}{\|\mathbf{TF}_{i,:}\|_{1}}$ ;
    // Clip document weights
// Step 3: Compute noisy document frequencies
$\mathbf{df}\leftarrow\sum_{i=1}^{n}\mathbf{TF}_{i,j}$ for $j\in|\mathcal{V}|$ ;
// Compute document frequencies
$\sigma\leftarrow\frac{\varepsilon_{\mathrm{idf}}}{w_{\max}+1}$ ;
// Noise scale
$\tilde{\mathbf{df}}\leftarrow\max(\mathbf{df}+\textsf{DLaplace}(\sigma),1)$ ;
// Add DP noise
$\tilde{n}\leftarrow n+\textsf{DLaplace}(\sigma)$ ;
// Noisy document count
// Step 4: Compute DP TF-IDF and extract keywords
[…]
// Compute DP IDF
$\mathbf{TFIDF}^{DP}\leftarrow\mathbf{TF}\odot\mathbf{idf}^{DP}$ ;
// Compute DP TF-IDF
$h_{i}\leftarrow$ top $k$ tokens from $\mathbf{TF}_{i,:}$ for each $i\in[n]$ ;
// Extract top $k$ keywords per conversation
// Step 5: Apply DP selection mechanism
$\mathcal{K}\leftarrow\mathsf{PPS}_{k,\varepsilon_{\mathrm{sel}},\delta_{\mathrm{sel}}}(h_{1},\ldots,h_{n})$ ;
// Select final keywords with DP
// Step 6: Limit to maximum keywords (optional)
if $|\mathcal{K}|>K$ then
 $\mathcal{K}\leftarrow\textsf{TopK}(\mathcal{K},K,\text{by original counts})$
// Step 7: LLM refinement
$\mathcal{K}\leftarrow\textsf{LLMRefineKeywords}(\mathcal{K},K)$ ;
// See prompt in §G.5
return $\mathcal{K}$
Algorithm 5 Differentially Private TF-IDF Keyword Extraction (KwSet-TFIDF)
Parameters: Maximum keywords per conversation $m$, maximum total keywords $K$.
Parameters: Privacy parameters $\varepsilon_{\mathrm{sel}}>0$, $\delta_{\mathrm{sel}}\in(0,1]$.
Input: Input dataset $\bm{x}=(x_{1},x_{2},\ldots,x_{n})\in\mathcal{X}^{n}$
Output: Differentially private keyword set $\mathcal{K}_{\text{refined}}$
1ex
// Step 1: Extract a set of at most $m$ keywords $h_{i}$ for each $x_{i}$.
for $i=1,\ldots,n$ do
 $h_{i}\leftarrow\mathsf{GetKeywords}(x_{i})$ ;
 // See prompt template in §G.5.
// Step 2: Apply DP selection mechanism
$\mathcal{K}\leftarrow\mathsf{PPS}_{k,\varepsilon_{\mathrm{sel}},\delta_{\mathrm{sel}}}(h_{1},\ldots,h_{n})$
// Step 3: LLM refinement
while $|\mathcal{K}|>K$ do
 $\mathcal{K}_{\text{refined}}\leftarrow\textsf{LLMQuery}(\textsf{RefinePrompt}(\mathcal{K}))$ ;
 // LLM refinement (see §G.5)
return $\mathcal{K}_{\text{refined}}$
Algorithm 6 DP Selection from Facet Keywords with LLM Refinement (KwSet-LLM)
Parameters: Initial batch size $n_{0}$, sequential batch size $b$, maximum initial keywords $K_{0}$.
Input: Input dataset $\mathbf{x}=(x_{1},x_{2},\ldots,x_{n})$ of structured facets
Output: Refined keyword set $\mathcal{K}^{pub}$
1ex
// Step 1: Extract text and split dataset
$\bm{x}_{\text{init}}\leftarrow\bm{x}[1:n_{0}]$, $\bm{x}_{\text{seq}}\leftarrow\bm{x}[n_{0}+1:n]$ ;
// Split into initial and sequential sets
// Step 2: Extract and refine initial keywords
$\mathbf{K}_{raw}\leftarrow\bigcup_{x\in\bm{x}_{init}}\{\textsf{NER}(x)\cup\textsf{NounChunks}(x)\cup\textsf{RAKE}(x)\}$ ;
// Multi-method extraction
$\mathcal{K}^{pub}\leftarrow\textsf{LLMRefine}(\mathbf{K}_{raw})$ ;
// Remove redundancy, merge concepts (see §G.6)
$\mathcal{K}^{pub}\leftarrow\textsf{RandomSample}(\mathcal{K}^{pub},\min(|\mathcal{K}^{pub}|,K_{0}))$ ;
// Limit initial set
// Step 3: Sequential updating with LLM feedback
for $i\leftarrow 1$ to $\lceil|\bm{x}_{seq}|/b\rceil$ do
 $\text{batch\_text}\leftarrow\textsf{Join}(\bm{x}_{seq}[(i-1)b+1:ib])$ ;
 // Merge batch conversations
 $(\mathbf{remove},\mathbf{append})\leftarrow\textsf{LLMUpdate}(\mathcal{K}^{pub},\text{batch\_text})$ ;
 // Get update suggestions (see §G.6)
 $\mathcal{K}^{pub}\leftarrow(\mathcal{K}^{pub}\setminus\mathbf{remove})\cup\mathbf{append}$ ;
 // Apply updates
return $\mathcal{K}^{pub}$
Algorithm 7 LLM-guided Sequential Keyword Refinement (KwSet-Public)
Parameters: Maximum number of keywords $k$ per conversation.
Parameters: Privacy parameters $\epsilon_{\mathrm{hyb}}>0$, $\delta_{\mathrm{hyb}}\in(0,1]$.
Input: Public keyword set $\mathcal{K}^{pub}$, input dataset $\bm{x}=(x_{1},x_{2},\ldots,x_{n})$
Output: Hybrid DP keyword set $\mathcal{K}^{hyb}$
1ex
// Step 1: LLM-guided candidate extraction in batches
for $i=1$ to $n$ do
 $h_{i}\leftarrow\textsf{GetNovelKeywords}(\mathcal{K}^{pub},x_{i})$ ;
 // Using prompt; see §G.7
// Step 2: Partition selection to select keywords corresponding to private conversations
$\mathcal{K}_{\text{private}}\leftarrow\mathsf{PPS}_{k,\varepsilon,\delta}(h_{1},\ldots,h_{n})$
return $\mathcal{K}_{\text{private}}\cup\mathcal{K}_{\text{pub}}$
Algorithm 8 Combined Public-Private DP Keyword Selection (KwSet-Hybrid)
Appendix D Summary Comparison Examples